OpenStudy (anonymous):

what is (1 over cos theta sin theta)cos theta in terms of sin theta?

OpenStudy (anonymous):

Is this how it looks? \[\frac{ 1 }{ \cos \Theta \sin \Theta }\sin \Theta\]

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

okay perfect! think about what can you cancel now?

OpenStudy (anonymous):

ahh my bad wait a sec

OpenStudy (anonymous):

\[\frac{ 1 }{ \cos \Theta \sin \Theta } \cos \Theta \]

OpenStudy (anonymous):

okay now think what you can cancel?

OpenStudy (anonymous):

the cos theta?

OpenStudy (anonymous):

yup! so what do you get?

OpenStudy (anonymous):

(1/sin theta)cos theta

OpenStudy (anonymous):

is that really it? thats all...?

OpenStudy (anonymous):

once you cancel the cos theta, there's no cos theta left

OpenStudy (anonymous):

you're only left with sin theta

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yeah, X) thanks so much!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Just a friendly reminder. The sin theta is in the denominator of the fraction, so 1 over sine theta is equivalent to cosecant theta, or \[\frac{ 1 }{\sin \Theta }=\csc \Theta \]

OpenStudy (anonymous):

right the ques however only asked in terms of sin

OpenStudy (anonymous):

oops, well, as long as he/she left it in the denominator. Many students tend to write the answer without the 1 on top.